OpenStudy (anonymous):

compute the derivative y=e^xlnx and y=(lnx+e^x)5

OpenStudy (anonymous):

1.\[y=e^xlnx\] 2.\[y=(lnx+e^x)^5 \]

OpenStudy (anonymous):

1º use derivative of the product rule: (uv)' = u'v+uv' let u=e^x and v = ln x so (e^xlnx)' =(e^x)'lnx+e^x(lnx)'=e^xlnx+e^x/x 2º use chain rule for derivative: [(lnx+e^x)^5]'=5(lnx+e^x)^4*(1/x +e^x)